# Unsolvable Problem Detection: Evaluating Trustworthiness of Large Multimodal Models Unsolvable Problem Detection: Robust Understanding Evaluation for Large Multimodal Models

### v2

@@ Abstract
Atsuyuki Miyai1 Jingkang Yang2 Jingyang Zhang3 Yifei Ming4 Qing Yu1,5 Go Irie6 Yixuan Li4 Hai Li3 Ziwei Liu2 Kiyoharu Aizawa1 1The University of Tokyo 2S-Lab, Nanyang Technological University 3Duke University 4University of Wisconsin-Madison 5LY Corporation 6Tokyo University of Science
@@ 1 Introduction
In recent years, following the revolutionary development of Large Language Models (LLMs) (Chen et al., 2024a; Chiang et al., 2023; Touvron et al., 2023; Wei et al., 2023), Large Multimodal Models (LMMs) (Liu et al., 2024c; Wang et al., 2023c; Hurst et al., 2024) have also demonstrated profound capabilities in various applications and significantly enhance the performance in image reasoning tasks (Antol et al., 2015; Liu et al., 2024b, e; Yue et al., 2024a).
Assessing the understanding capability of LMMs is crucial for advancing fundamental progress. Multiple-Choice Question Answering (MCQA) serves as a fundamental format for understanding evaluation and is widely used in well-established benchmarks such as MMBench Liu et al. (2024e) and MMMU Yue et al. (2024a). Each MCQA instance consists of a question paired with multiple answer options, requiring models to select the correct one. MCQA enables precise evaluation of LMMs and facilitates solid progress in the field. Consequently, many MCQA-based benchmarks have been proposed recently (Fu et al., 2024; Yue et al., 2024b; Hu et al., 2025).
@@ 5.2 Main Results
First, we describe the two most crucial findings (F1 and F2 below).
@@
However, among open-source LMMs, models with large LLMs such as LLaVA-NeXT-34B and InternVL2-40B demonstrate performance comparable to closed-source models. Compared to smaller models trained on the same VQA data, such as LLaVA-NeXT-13B and InternVL2-2B/8B, there is a significant performance improvement, suggesting that the performance of the base LLM also plays a crucial role.
@@
F3: UPD Score is Significantly Lower than Standard in Base and Solution Varies by LMMs.
Fig. 3 shows the Standard (blue) and UPD (red) accuracy. The performance was compared, with each row showing the results for AAD, IASD, and IVQD, and each column showing the results for Base, Option, and Instruction. Models (i)-(v) in the figure denotes open-source models and Model (vi)-(viii) denotes closed-source models.
@@
F4: Performance of AAD, IASD, and IVQD Diagnose Each LMM’s Weakness.
@@
F5: Performance Trends Vary across Abilities.
@@ Appendix A Additional Related Work
AI Safety. A reliable visual recognition system should not only produce accurate predictions on known context but also detect unknown examples (Amodei et al., 2016; Mohseni et al., 2022; Hendrycks et al., 2021; Hendrycks and Mazeika, 2022). The representative research field to address this safety aspect is out-of-distribution (OOD) detection (Hendrycks and Gimpel, 2017; Liang et al., 2018; Yang et al., 2024b, 2022; Zhang et al., 2024a). OOD detection is the task of detecting unknown samples during inference to ensure the safety of the in-distribution (ID) classifiers. Along with the evolution of the close-set classifiers, the target tasks for OOD detection have evolved from the detectors for conventional single-modal classifiers to recent CLIP-based methods (Miyai et al., 2024; Hendrycks and Gimpel, 2017; Yu and Aizawa, 2019; Wang et al., 2021; Du et al., 2022; Ming et al., 2022b; Esmaeilpour et al., 2022; Ming et al., 2022a; Yang et al., 2023; Wang et al., 2023a; Miyai et al., 2023a, b). The next crucial challenge is to evolve the problems faced in OOD detection to LMMs in the VQA task. We consider that our UPD is an extension of the concept of OOD detection, where the model should detect and not predict unexpected input data.
@@ Appendix B Benchmark Construction
B.1 Processing for MMBench Adaptation
B.2 Construction of MM-AAD Bench
B.3 Construction of MM-IASD Bench
@@ B.4 Construction of MM-IVQD Bench
To create MM-IVQD Bench, we first exclude the questions that can be relevant to most images and then shuffle the original image-question pairs. In Table C, we show some representative examples of removed questions. For example, the question of “How many …" can be compatible with any image, since the correct option of “None of the above" always exists for any image even when the image has no corresponding objects. For the question of “What’s the profession …", we can interpret the profession from any kind of image (e.g., A beautifully captured image would suggest the profession of a photographer). In addition, we exclude the option “Can’t judge from the image due to the lack of image information.” because this option can be a correct answer for IVQD questions. Again, we conduct a manual check to guarantee the incompatibility of image-question pairs. Our MM-IVQD Bench has 356 IVQD questions over 12 abilities. The distribution of questions for each ability is shown in the bottom of Table B. Here, the lack of some ability (e.g.,#16 Image Emotion) indicates that there are many removed questions that can be applied to any image. Note that the small number of IVQD questions compared to AAD and IASD is due to our careful annotation and that even this number of questions is sufficient to show the performance difference between each LMM and method from our main experimental results.
@@
B.5 Manual Curation Procedure
@@
B.6 Validity of UPD Benchmark on More Complex Datasets
@@
Result. We show the comparison results in Table D. Based on these results, in contrast to MM-UPD, we could not verify the efficacy of either the Option or Instruction approaches. This result reveals that the evaluation using MMMU fails to capture important findings of the effectiveness of these prompting approaches for UPD. Specifically, for expert-level problems, LMMs do not have accurate answers due to the lack of capability. Therefore, even if they choose an incorrect option when encountering an unsolvable problem, this only indicates a lack of reasoning ability or knowledge and does not necessarily demonstrate a lack of refusal ability. Additionally, due to the very low overall performance, it becomes difficult to have meaningful discussions based on these minute differences in scores. Therefore, we exclude datasets with low Standard accuracy.
@@ C.2 Detail of LLM-driven Methods
Self-reflection Self-reflection is a method that allows the model to reflect on its own responses (Kadavath et al., 2022). It has been shown that LLMs might have preliminary capabilities for judging and evaluating their own answers (Kadavath et al., 2022; Feng et al., 2024). In this experiment, we evaluate whether self-reflection is effective for UPD. We show the prompt for self-reflection in Table F. We prompt the LMM to self-reflect directly after its generated answer with the phrase “The above answer is: 1. True 2. False,” following LLM protocols (Kadavath et al., 2022; Feng et al., 2024). For evaluation, if the LMM outputs “2. False," the response will be withdrawn. Otherwise, we use the original LMM’s response for the evaluation.
@@ D.1 Setup
Dataset. For the dataset, we use a subset of an open-knowledge VQA dataset, A-OKVQA (Schwenk et al., 2022). It is a single-choice type VQA dataset that has been used for training InstructBLIP (Dai et al., 2023) and LLaVA-1.5 (Liu et al., 2024c). The samples in A-OKVQA do not overlap with our benchmarks.
@@ D.3 Result
Table E demonstrates that instruction tuning is effective for UPD, showing the performance efficacy and limitations with UPD-specific training. However, UPD-specific training may degrade the performance of other general tasks. Therefore, if the user intends to use LMMs for broader, more general purposes rather than just for UPD tasks, instruction tuning may not be a good approach. It is a future challenge to propose a method that improves UPD performance while maintaining performance on general tasks.

### v4

@@ Abstract
Atsuyuki Miyai1 Jingkang Yang2 Jingyang Zhang3 Yifei Ming4 Qing Yu1,5 Go Irie6 Yixuan Li4 Hai Li3 Ziwei Liu2 Kiyoharu Aizawa1,6 1The University of Tokyo 2S-Lab, Nanyang Technological University 3Duke University 4University of Wisconsin-Madison 5LY Corporation 6Tokyo University of Science miyai@cvm.t.u-tokyo.ac.jp
@@ 1 Introduction
In recent years, following the revolutionary development of Large Language Models (LLMs) (Chen et al., 2024a; Chiang et al., 2023; Touvron et al., 2023; Wei et al., 2023), Large Multimodal Models (LMMs, also referred to as Multimodal Large Language Models or MLLMs) (Liu et al., 2024c; Wang et al., 2023c; Hurst et al., 2024) have also demonstrated profound capabilities in various applications and significantly enhance the performance in image reasoning tasks (Antol et al., 2015; Liu et al., 2024b, e; Yue et al., 2024a).
Assessing the understanding capability of LMMs is crucial for advancing fundamental progress. Multiple-Choice Question Answering (MCQA) serves as a fundamental format for understanding evaluation and is widely used in well-established benchmarks such as MMBench Liu et al. (2024e) and MMMU Yue et al. (2024a). Each MCQA instance consists of a question paired with multiple answer options, requiring models to select the correct one. MCQA enables precise evaluation of LMMs and facilitates solid progress in the field. Consequently, many MCQA-based benchmarks have been proposed recently (Fu et al., 2024; Yue et al., 2024b; Hu et al., 2025; Onohara et al., 2025).
@@ 5.2 Main Results
First, we describe the three most crucial findings (F1, F2, and F3 below).
@@
F3. Larger Open-source LMMs Mitigate the Gap.
Among open-source LMMs, models with large LLMs such as LLaVA-NeXT-34B and InternVL2-40B demonstrate performance comparable to closed-source models. Compared to smaller models trained on the same VQA data, such as LLaVA-NeXT-13B and InternVL2-2B/8B, there is a significant performance improvement, suggesting that the performance of the base LLM also plays a crucial role.
@@
F4: UPD Score is Significantly Lower than Standard in Base and Solution Varies by LMMs.
Fig. 3 shows the Standard (blue) and UPD (red) accuracy. The performance was compared, with each row showing the results for AAD, IASD, and IVQD, and each column showing the results for Base, Option, and Instruction. Models (i)-(v) in the figure denote open-source models and Models (vi)-(viii) denote closed-source models.
@@
F5: Performance of AAD, IASD, and IVQD Diagnose Each LMM’s Weakness.
@@
F6: Performance Trends Vary across Abilities.
@@ 5.4 Bottleneck Analysis
5.5 The Effect of Fine-tuning
We provide a brief discussion on the performance gains obtained through fine-tuning on UPD-like data. For the training dataset, we use a subset of an open-knowledge VQA dataset, A-OKVQA (Schwenk et al., 2022). The samples in A-OKVQA do not overlap with our benchmarks.
Our training results suggest the following key observations: (i) The performance is sensitive to the composition ratio of Standard, AAD, and IVQD samples in the training set. The optimal recipe was found to be a ratio of 0.6 for Standard, 0.2 for AAD, and 0.2 for IVQD, while excluding IASD entirely. (ii) Compared to the prompt-based approach, fine-tuning on UPD-like data yields measurable performance improvements. However, we note that specializing the model for UPD may lead to degraded performance on general-purpose tasks, indicating that this strategy may not represent a universally optimal solution.
Further details and experimental results are provided in Appendix D.
@@ 6 Conclusion
Acknowledgment
This work was supported by JSPS 25H01164 and JST BOOST, Japan Grant Number JPMJBS2418.
@@ Appendix A Additional Related Work
AI Safety. A reliable visual recognition system should not only produce accurate predictions on known context but also detect unknown examples (Amodei et al., 2016; Mohseni et al., 2022; Hendrycks et al., 2021; Hendrycks and Mazeika, 2022). The representative research field to address this safety aspect is out-of-distribution (OOD) detection (Hendrycks and Gimpel, 2017; Liang et al., 2018; Yang et al., 2024b, 2022; Zhang et al., 2024a). OOD detection is the task of detecting unknown samples during inference to ensure the safety of the in-distribution (ID) classifiers. Along with the evolution of the close-set classifiers, the target tasks for OOD detection have evolved from the detectors for conventional single-modal classifiers to recent CLIP-based methods (Miyai et al., 2024; Hendrycks and Gimpel, 2017; Yu and Aizawa, 2019; Wang et al., 2021; Du et al., 2022; Ming et al., 2022b; Esmaeilpour et al., 2022; Ming et al., 2022a; Yang et al., 2023; Wang et al., 2023a; Miyai et al., 2023a, b; Noda et al., 2025). The next crucial challenge is to evolve the problems faced in OOD detection to LMMs in the VQA task. We consider that our UPD is an extension of the concept of OOD detection, where the model should detect and not predict unexpected input data.
@@ Appendix B Benchmark Construction
B.1 Processing for MMBench Adaptation
| Model | AAD (Base) | IASD (Base) | IVQD (Base) |
| InternVL2-8B | 39.83 ± 1.51 (41.7, 39.8, 38.0) | 48.03 ± 0.97 (49.4, 47.4, 47.3) | 37.37 ± 0.60 (37.1, 38.2, 36.8) |
| LLaVA-OV-7B | 7.93 ± 0.29 (7.9, 8.3, 7.6) | 8.83 ± 0.40 (8.5, 8.6, 9.4) | 3.67 ± 0.38 (4.2, 3.4, 3.4) |
| InternVL2-40B | 39.80 ± 1.80 (42.1, 39.6, 37.7) | 47.77 ± 0.52 (48.5, 47.4, 47.4) | 37.53 ± 0.74 (36.7, 38.5, 37.4) |
| GPT-4o | 55.37 ± 1.06 (54.1, 56.7, 55.3) | 68.40 ± 0.64 (68.9, 67.5, 68.8) | 70.60 ± 1.08 (71.6, 69.1, 71.1) |
B.2 Construction of MM-AAD Bench
B.3 Construction of MM-IASD Bench
@@ B.4 Construction of MM-IVQD Bench
To create MM-IVQD Bench, we first exclude the questions that can be relevant to most images and then shuffle the original image-question pairs. In Table C, we show some representative examples of removed questions. For example, the question of “How many …" can be compatible with any image, since the correct option of “None of the above" always exists for any image even when the image has no corresponding objects. For the question of “What’s the profession …", we can interpret the profession from any kind of image (e.g., A beautifully captured image would suggest the profession of a photographer). In addition, we exclude the option “Can’t judge from the image due to the lack of image information.” because this option can be a correct answer for IVQD questions. Again, we conduct a manual check to guarantee the incompatibility of image-question pairs. Our MM-IVQD Bench has 356 IVQD questions over 12 abilities. The distribution of questions for each ability is shown in the bottom of Table B. Here, the lack of some ability (e.g.,#16 Image Emotion) indicates that there are many removed questions that can be applied to any image. Note that the small number of IVQD questions compared to AAD and IASD is due to our careful annotation. The additional experiments in Sec. B.5 indicate even this number of questions is sufficient to show the performance difference between each LMM and method from our main experimental results.
@@
B.5 Performance Variance on IVQD
We demonstrate that the dataset size is sufficient by showing that the performance variance remains small under different conditions, such as shifting the positions of answer options. We conducted additional experiments using three different patterns based on option shifting and measured the accuracy for each. For unsolvable problems with only two answer choices, a third shift pattern does not exist. In such cases, we reused the questions from Pattern 2 for Pattern 3. The proportion of two-choice questions is 10.1% in AAD, 1.85% in IASD, and 8.7% in IVQD.
We show the results in Table D. The results show that the variance in IVQD is similarly small compared to AAD and IASD, which supports the reliability of the evaluation in terms of dataset size.
B.6 Manual Curation Procedure
@@
B.7 Validity of UPD Benchmark on More Complex Datasets
@@
Result. We show the comparison results in Table E. Based on these results, in contrast to MM-UPD, we could not verify the efficacy of either the Option or Instruction approaches. This result reveals that the evaluation using MMMU fails to capture important findings of the effectiveness of these prompting approaches for UPD. Specifically, for expert-level problems, LMMs do not have accurate answers due to the lack of capability. Therefore, even if they choose an incorrect option when encountering an unsolvable problem, this only indicates a lack of reasoning ability or knowledge and does not necessarily demonstrate a lack of refusal ability. Additionally, due to the very low overall performance, it becomes difficult to have meaningful discussions based on these minute differences in scores. Therefore, we exclude datasets with low Standard accuracy.
@@ C.2 Detail of LLM-driven Methods
Self-reflection Self-reflection is a method that allows the model to reflect on its own responses (Kadavath et al., 2022). It has been shown that LLMs might have preliminary capabilities for judging and evaluating their own answers (Kadavath et al., 2022; Feng et al., 2024). In this experiment, we evaluate whether self-reflection is effective for UPD. We show the prompt for self-reflection in Table G. We prompt the LMM to self-reflect directly after its generated answer with the phrase “The above answer is: 1. True 2. False,” following LLM protocols (Kadavath et al., 2022; Feng et al., 2024). For evaluation, if the LMM outputs “2. False," the response will be withdrawn. Otherwise, we use the original LMM’s response for the evaluation.
@@ D.1 Setup
Dataset. For the dataset, we use a subset of an open-knowledge VQA dataset, A-OKVQA (Schwenk et al., 2022). It is a multiple-choice type VQA dataset that has been used for training InstructBLIP (Dai et al., 2023) and LLaVA-1.5 (Liu et al., 2024c). The samples in A-OKVQA do not overlap with our benchmarks.
@@ D.3 Result
Table F demonstrates that instruction tuning is effective for UPD, showing the performance efficacy and limitations with UPD-specific training. However, UPD-specific training may degrade the performance of other general tasks. Therefore, if the user intends to use LMMs for broader, more general purposes rather than just for UPD tasks, instruction tuning may not be a good approach. It is a future challenge to propose a method that improves UPD performance while maintaining performance on general tasks.